OpenStudy (anonymous):

A wheel is marked off into 8 equal sections numbered 1 through 8. 1,4 and 7 are colored red. You spin the wheel... What is the probability the outcome will be either red OR and even number? a) 4/8 b) 3/8 c) 7/8 d) 6/8

OpenStudy (anonymous):

I got 3/16...

OpenStudy (l):

Red sections 1 4 7 Even Sections 2 4 6 8 So red or even sections 1 2 4 6 7 8 So the probability is 6 / 8 = 3 / 4 = 0.75 (d) *

OpenStudy (anonymous):

Thank you